# Learning Enhanced Ensemble Filters

### v1

@@ 4 Particle Approximation of Learned Mean-field Filters
After the prediction steps (4.1a) and (4.1b) we obtain a set of particles $\{(\widehat{v}^{(n)}_{j+1},h(\widehat{v}^{(n)}_{j+1}))\}_{n=1}^{N}$. The computation of $K_{\theta}$ is elaborated through equations (2.22), (2.23), and (2.24) in a mean-field perspective, now extended to also include this particle setting. As in Section 2, we drop time index $j$ in what follows and recall that $\rho_{h}=\operatorname{Law}\bigl{(}(\widehat{v},h(\widehat{v}))\bigr{)}$ is the joint distribution of $(\widehat{v},h(\widehat{v}))$ in the mean-field limit. In the particle case, under their implied empirical distribution, we have
@@
and use this as input measure to $K_{\theta}.$ The core idea is to train a parameterized neural operator $\mathfrak{F}\left(\widehat{v},h(\widehat{v}),y^{\dagger},\rho_{h};\theta\right)$ that outputs the correction terms used to compute $K_{\theta}$. This can be done using the transformer measure neural mapping architecture $\mathsf{F}$ (3.3) introduced in Section 3. When applied to empirical measures such as (4.2), this scheme is equivalent to the set transformer architecture [44] acting on sequences, as outlined in Remark 5. In this section, which is concerned with implementation, we will work with the formulation of the scheme on sequences; we will also show that the resulting scheme is invaraint to permutation of the elements of the sequence. To distinguish the resulting operator on sequences to the one acting on measures from Section 3 we will employ the notation $\mathsf{F}^{\textrm{ST}}$ opposed to $\mathsf{F}$; see (4.18) for the definition of $\mathsf{F}^{\textrm{ST}}$. Due to the permutation invariance property, we can treat the input as an unordered set: we define the corresponding operator acting on finite sets using the notation $F^{\mathrm{ST}}$ ; see (4.25).
The parameters $\theta$ learned in $K_{\theta}$ implicitly depend on $N$, the number of ensemble members used in training. However, the relation between the set transformer and its mean field limit from Section 3 allows us to deploy the model parameters $\theta$, trained using an ensemble size $N$, to an ensemble of a different sizes $N^{\prime}$. This is because the $N$ and $N^{\prime}$ systems may both be thought of as approximating the same mean-field limit. Similarly to standard use of the EnKF, our learning framework also incorporates techniques such as inflation and localization to enhance the performance of the filter [6] as ensemble sizes vary. We view this as a fine-tuning: fixing the majority of the parameters in $\theta$ as ensemble sizes vary, but allowing a small subset of parameters—including those related to inflation and localization—to vary with ensemble size.
We present the complete implementation details of our method in the following subsections. In Subsection 4.1, we review the set transformer architecture. In Subsection 4.2, we present our architecture that learns a parameterized gain matrix, generalizing the EnKF formulation, and enhanced by use of adaptively learned inflation and localization parameters. Subsection 4.3 explains the loss function and training process. In Subsection 4.4, we describe how to efficiently fine-tune our method to allow for ensemble size dependent inflation and localization. In A.1 we provide a comprehensive summary of our learning framework with detailed pseudo-code algorithms.
@@ 4.1 The Set Transformer
In the context of practical implementations, $\mathsf{A}^{\textrm{ST}}$ is commonly instantiated as a multihead attention mechanism, which allows the model to jointly attend to information from different representation subspaces. The detailed formulation of multihead attention is presented in A.2.
@@
Here $\mathsf{A}^{\textrm{ST}}$ is the sequence-to-sequence attention (4.5); $\mathsf{F}^{\textrm{LN}}$ denotes application of the layer normalization $\mathfrak{F}^{\textrm{LN}}$ (3.6), elementwise in the sequence; $\mathsf{F}^{\textrm{NN}}$ denotes application of the feedforward operator $\mathfrak{F}^{\textrm{NN}}$ (3.8), also elementwise in the sequence. The $+$ operator acting on sequences is interpreted as adding elements with the same index, i.e.,
@@
where the layer normalization $\mathsf{F}^{\textrm{LN}}_{1},\mathsf{F}^{\textrm{LN}}_{2}$ have trainable parameters according to (3.6)(3.7), the $\mathsf{F}^{\textrm{NN}}$ has trainable parameters according to (3.8), and the attention layer $\mathsf{A}^{\textrm{ST}}$ has trainable parameters according to (4.7). We note that practically the cross-attention block on sequences is the multi-head attention block from [32, 44] and may be viewed as the sequence-to-sequence analog of the cross-attention block $\mathsf{C}$ on measures from (3.4b).
The set transformer architecture is composed of several self-attention blocks (SAB) (Definition 7) and a pooling by multi-head attention block (PMA) (Definition 8), which are defined using CAB.
@@
The pooling by multi-head attention block (PMA) $\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ is defined by replacing the first input in the CAB $\mathsf{C}^{\textrm{ST}}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ by a trainable parameter $s$ called the seed. The seed $s\in\mathcal{U}([N_{s}],\mathbb{R}^{d_{s}})$ is a sequence, where $N_{s}$ and $d_{s}$ are fixed hyperparameters. Therefore $\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ maps a sequence of any finite length to a sequence with fixed length,
@@
The subscripts $\ell\in[N_{\mathrm{e}}+N_{\mathrm{d}}]$ distinguish the dependence of SABs on different trainable parameters. Integer $N_{\mathrm{e}}$ is the number of SABs, $\mathsf{S}^{\textrm{ST}}_{\ell},\ell=1,2,\ldots,N_{\mathrm{e}}$, in the encoder $\mathsf{F}^{\textrm{Enc}}$, while $N_{\mathrm{d}}$ is the number of SABs, $\mathsf{S}^{\textrm{ST}}_{\ell},\ell=1+N_{\mathrm{e}},2+N_{\mathrm{e}},\ldots,N_{\mathrm{d}}+N_{\mathrm{e}}$, in the decoder $\mathsf{F}^{\textrm{Dec}}$. Specifically, we have
@@
Further details about the our practical set transformer architecture, including the dimensions of latent features and the choices of layers, are provided in B.2.
@@
$\begin{split}&\theta(\mathsf{F}^{\textrm{ST}})=\theta\bigl{(}\mathsf{F}^{\textrm{Dec}}\bigr{)}\cup\theta\bigl{(}\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{)}\cup\theta\bigl{(}\mathsf{F}^{\textrm{Enc}}\bigr{)}\\ &=\theta\bigl{(}\mathsf{F}^{\textrm{NN}}_{\mathrm{in}}\bigr{)}\cup\theta\bigl{(}\mathsf{F}^{\textrm{NN}}_{\mathrm{out}}\bigr{)}\cup\bigl{(}\cup_{\ell=1}^{N_{\mathrm{e}}+N_{\mathrm{d}}}\theta\bigl{(}\mathsf{S}^{\textrm{ST}}_{\ell}\bigr{)}\bigr{)}\cup\theta\bigl{(}\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{)}.\end{split}$ (4.20)
The set transformer architecture exhibits two key properties when processing an input sequence, namely, the permutation invariance (Proposition 2) and adaptation to variant input lengths (Proposition 3). These two propositions are proved in [44] under a slightly different setting: in the original work a length-$N$ sequence in $\mathbb{R}^{d}$ is represented as an $N\times d$ matrix, whereas here it is a mapping $u\in\mathcal{U}([N],\mathbb{R}^{d})$.
@@ 4.2.1 Learning the Gain Matrix
Recall that, in the mean-field limit, we impose an inductive bias on our archiecture by defining $K_{\theta}^{(1)}$ and $K_{\theta}^{(2)}$ in the form (2.23), allowing for a learned correction through (2.24). The empirical version of (2.23), used in the practical implementation, takes the form
@@
where $\widehat{w}^{(n)}_{\theta}$ and $\widehat{z}^{(n)}_{\theta}$ are trainable correction terms.
The correction terms are given by a neural operator $\mathfrak{F}$ in the mean-field perspective (2.24) (Section 2). In practical implementation, we train a neural network $F^{\mathrm{gain}}$ incorporating a set transformer architecture to output correction terms for each particle. We replace the measure dependencies $\rho_{h}$ in $\mathfrak{F}$ by the ensemble $\{(\widehat{v}^{(\ell)},h(\widehat{v}^{(\ell)}))\}_{\ell=1}^{N}$ and use the set transformer $F^{\mathrm{ST}}$ (4.25) to process the ensemble into a feature vector $f_{v}$. Recall that this can be viewed as using the empirical approximation $\rho_{h}^{(N)}\approx\rho_{h}$ given by (4.2).
We consider a neural network $F^{\mathrm{gain}}:\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{\text{ST}}}\rightarrow\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}$ with the standard multilayer perceptron (MLP) architecture to learn the correction terms:
@@
From (4.27a) and (4.27b), it is evident that if the correction terms $\widehat{w}^{(n)}_{\theta}=\widehat{z}^{(n)}_{\theta}=0$, the formulation reduces to the EnKF (2.16). When the dynamics and observation models are linear, or in other words, when the filtering distribution of the states $v$ is Gaussian, the Kalman filter achieves optimality [7, 8, 40]. Therefore, the motivation for our proposed scheme (2.22) lies in its optimality in the linear (or Gaussian) setting, while further aiming to improve the performance of the Kalman filter in nonlinear problems by learning the correction terms.
@@
We have implemented the direct end-to-end learning of the gain matrix $K_{\theta}$ in addition to the proposed approach with correction terms. We find that the end-to-end approach results in unstable and slower learning. A poorly initialized or an unconstrained $K_{\theta}$ leads to substantial divergence from the ground truth. The EnKF-like structure in (4.26b) with learnable correction terms provides better stability while maintaining adaptability to nonlinear dynamics.
@@ 4.2.2 Learning the Inflation
The most common form of inflation is multiplicative inflation, where the analysis ensemble $\{v^{(n)}\}_{n=1}^{N}$ calculated after the analysis step (4.1c) is modified as follows
@@
where $\alpha\geq 1$ is the inflation parameter. In our learning framework, we replace the term proportional to $(\alpha-1)$ by a learned correction term $\widehat{u}^{(n)}_{\theta}$:
$\widehat{v}^{(n)}\leftarrow\widehat{v}^{(n)}+\widehat{u}^{(n)}_{\theta}.$ (4.30)
Since the inflation is to increase the ensemble spread, and by comparison with (4.29), we consider an MLP-based neural network $F^{\mathrm{infl}}:\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{\text{ST}}}\rightarrow\mathbb{R}^{d_{v}}$ given by
$F^{\mathrm{infl}}\bigl{(}\widehat{v}^{(n)},f_{v};\theta_{\mathrm{infl}}\bigr{)}=\widehat{u}^{(n)}_{\theta},$ (4.31)
where $\theta_{\mathrm{infl}}$ denotes trainable parameters. Recall that the feature vector $f_{v}$ is computed solely from the ensemble $\{(\widehat{v}^{(n)},h(\widehat{v}^{(n)}))\}_{n=1}^{N}$, according to (4.25). We substitute (4.30) into our parameterized analysis step (4.1c), writing the new analysis step with incorporated inflation in the form
$v^{(n)}=(\widehat{v}^{(n)}+\widehat{u}^{(n)}_{\theta})+K_{\theta}\left(y^{\dagger}-\widehat{y}^{(n)}\right).$ (4.32)
@@
The learned inflation term $\widehat{u}^{(n)}_{\theta}$ in (4.30) depends on the ensemble states $\widehat{v}^{(n)}$ and the feature vector $f_{v}$. The inflation term $\widehat{u}_{\theta}$ does not include information about the true observation $y^{\dagger}$ or the observation noise covariance $\Gamma$. Therefore, we avoid the risk that the inflation term would evolve into an end-to-end correction term, ensuring the learning performed by other components in our framework remains meaningful and effective.

### v2

@@ 4 Particle Approximation of Learned Mean-field Filters
After the prediction steps (4.1a) and (4.1b), we obtain a set of particles $\{(\widehat{v}^{(n)}_{j+1},h(\widehat{v}^{(n)}_{j+1}))\}_{n=1}^{N}$. The computation of $K_{\theta}$ is elaborated through equations (2.22), (2.23), and (2.24) in a mean-field perspective, now extended to also include this particle setting. As in Section 2, we drop time index $j$ in what follows and recall that $\rho_{h}=\operatorname{Law}\bigl{(}(\widehat{v},h(\widehat{v}))\bigr{)}$ is the joint distribution of $(\widehat{v},h(\widehat{v}))$ in the mean-field limit. In the particle case, under their implied empirical distribution, we have
@@
and use this as an input measure to $K_{\theta}.$ The core idea is to train a parameterized neural operator $\mathfrak{F}\left(\widehat{v},h(\widehat{v}),y^{\dagger},\rho_{h};\theta\right)$ that outputs the correction terms used to compute $K_{\theta}$. This can be done using the transformer measure neural mapping (TMNM) architecture $\mathsf{F}$ (3.3) introduced in Section 3. When applied to empirical measures such as (4.2), this TMNM scheme is equivalent to the set transformer architecture [44] acting on sequences, as outlined in Remark 5. In this section, which is concerned with implementation, we will work with the formulation of the scheme on sequences; we will also show that the resulting scheme is invariant to permutation of the elements of the sequence. To distinguish the resulting operator on sequences to the one acting on measures from Section 3, we will employ the notation $\mathsf{F}^{\textrm{ST}}$ opposed to $\mathsf{F}$; see (4.18) for the definition of $\mathsf{F}^{\textrm{ST}}$. Due to the permutation invariance property of the set transformer, we can treat the input as an unordered set: we define the corresponding operator acting on finite sets using the notation $F^{\mathrm{ST}}$ ; see (4.25).
The parameters $\theta$ learned in $K_{\theta}$ implicitly depend on $N$, the number of ensemble members used in training. However, the relation between the set transformer and its mean-field limit from Section 3 allows us to deploy the model parameters $\theta$, trained using an ensemble size $N$, to an ensemble of a different sizes $N^{\prime}$. This is because the $N$ and $N^{\prime}$ systems may both be thought of as approximating the same mean-field limit. Similarly to the standard use of the EnKF, our learning framework also incorporates techniques such as inflation and localization to enhance the performance of the filter [6] as ensemble sizes vary. We view this as a fine-tuning: fixing the majority of the parameters in $\theta$ as ensemble sizes vary, but allowing a small subset of parameters—including those related to inflation and localization—to vary with ensemble size.
We present the complete implementation details of our method in the following subsections. In Subsection 4.1, we review the set transformer architecture. In Subsection 4.2, we present our architecture that learns a parameterized gain matrix, generalizing the EnKF formulation, and enhanced by use of adaptively learned inflation and localization parameters. Subsection 4.3 explains the loss function and training process. In Subsection 4.4, we describe how to efficiently fine-tune our method to allow for ensemble size dependent inflation and localization. In A we provide a comprehensive summary of our learning framework with detailed pseudo-code algorithms.
@@ 4.1 The Set Transformer
In the context of practical implementations, $\mathsf{A}^{\textrm{ST}}$ is commonly instantiated as a multihead attention mechanism, which allows the model to jointly attend to information from different representation subspaces [32]. The detailed formulation of multihead attention is presented in B.
@@
Here $\mathsf{A}^{\textrm{ST}}$ is the sequence-to-sequence attention (4.5); $\mathsf{F}^{\textrm{LN}}$ denotes application of the layer normalization operator $\mathfrak{F}^{\textrm{LN}}$ (3.6), elementwise in the sequence; $\mathsf{F}^{\textrm{NN}}$ denotes application of the feedforward operator $\mathfrak{F}^{\textrm{NN}}$ (3.8), also elementwise in the sequence. The $+$ operator acting on sequences is interpreted as adding elements with the same index, i.e.,
@@
where the layer normalization layers $\mathsf{F}^{\textrm{LN}}_{1},\mathsf{F}^{\textrm{LN}}_{2}$ have trainable parameters according to (3.6)(3.7), the feedforward layer $\mathsf{F}^{\textrm{NN}}$ has trainable parameters according to (3.8), and the attention layer $\mathsf{A}^{\textrm{ST}}$ has trainable parameters according to (4.7). We note that practically the cross-attention block on sequences is the multi-head attention block from [32, 44] and may be viewed as the sequence-to-sequence analog of the cross-attention block $\mathsf{C}$ on measures from (3.4b).
The set transformer architecture is composed of several self-attention blocks (SAB) (Definition 7) and a pooling by multi-head attention block (PMA) (Definition 8), which are defined in the following using CAB.
@@
The pooling by multi-head attention block (PMA) $\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ is defined by replacing the first input in the CAB $\mathsf{C}^{\textrm{ST}}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ by a trainable parameter $s$, called the seed. The seed $s\in\mathcal{U}([N_{s}],\mathbb{R}^{d_{s}})$ is a sequence, where $N_{s}$ and $d_{s}$ are fixed hyperparameters. Therefore $\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ maps a sequence of any finite length to a sequence with fixed length,
@@
The subscripts $\ell\in[N_{\mathrm{e}}+N_{\mathrm{d}}]$ emphasize that the trainable parameters of these SABs are different. Integer $N_{\mathrm{e}}$ is the number of SABs, $\mathsf{S}^{\textrm{ST}}_{\ell},\ell=1,2,\ldots,N_{\mathrm{e}}$, in the encoder $\mathsf{F}^{\textrm{Enc}}$, while $N_{\mathrm{d}}$ is the number of SABs, $\mathsf{S}^{\textrm{ST}}_{\ell},\ell=1+N_{\mathrm{e}},2+N_{\mathrm{e}},\ldots,N_{\mathrm{d}}+N_{\mathrm{e}}$, in the decoder $\mathsf{F}^{\textrm{Dec}}$. Specifically, we have
@@
Further practical details about our set transformer architecture, including the dimensions of latent features and the choices of layers, are provided in C.1.
@@
$\begin{split}\theta(\mathsf{F}^{\textrm{ST}})&=\theta\bigl{(}\mathsf{F}^{\textrm{Dec}}\bigr{)}\cup\theta\bigl{(}\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{)}\cup\theta\bigl{(}\mathsf{F}^{\textrm{Enc}}\bigr{)}\\ &=\theta\bigl{(}\mathsf{F}^{\textrm{NN}}_{\mathrm{in}}\bigr{)}\cup\theta\bigl{(}\mathsf{F}^{\textrm{NN}}_{\mathrm{out}}\bigr{)}\cup\bigl{(}\cup_{\ell=1}^{N_{\mathrm{e}}+N_{\mathrm{d}}}\theta\bigl{(}\mathsf{S}^{\textrm{ST}}_{\ell}\bigr{)}\bigr{)}\cup\theta\bigl{(}\mathsf{C}^{\textrm{ST}}(s,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{)}.\end{split}$ (4.20)
The set transformer architecture exhibits two key properties when processing an input sequence, namely, the permutation invariance (Proposition 2) and adaptation to variable input lengths (Proposition 3). These two propositions are proved in [44] under a slightly different setting: in the original work a length-$N$ sequence in $\mathbb{R}^{d}$ is represented as an $N\times d$ matrix, whereas here it is a mapping $u\in\mathcal{U}([N],\mathbb{R}^{d})$.
@@ 4.2.1 Learning the Gain Matrix
Recall that, in the mean-field limit, we impose an inductive bias on our archiecture by defining $K_{\theta}^{(1)}$ and $K_{\theta}^{(2)}$ in the form (2.23), allowing for a learned correction through (2.24). Recall that $\widehat{m}$ and $\widehat{h}$ are the mean of $\{\widehat{v}^{(n)}\}_{n=1}^{N}$ and $\{h(\widehat{v}^{(n)})\}_{n=1}^{N}$, respectively. The empirical version of (2.23), used in the practical implementation, takes the form
@@
where $\widehat{w}^{(n)}_{\theta}$ and $\widehat{z}^{(n)}_{\theta}$ are trainable correction terms that depend on the state $\widehat{v}^{(n)}$, the observation $h(\widehat{v}^{(n)})$, the true observation $y^{\dagger}$, and the ensemble $\{(\widehat{v}^{(n)},h(\widehat{v}^{(n)}))\}_{n=1}^{N}$.
The correction terms are defined by a neural operator $\mathfrak{F}$ as in (2.24) for the mean-field perspective presented in Section 2. In our practical implementation, we replace the dependence on $\rho_{h}$ in $\mathfrak{F}$ by the ensemble $\{(\widehat{v}^{(\ell)},h(\widehat{v}^{(\ell)}))\}_{\ell=1}^{N}$ and use the set transformer $F^{\mathrm{ST}}$ (4.25) to process the ensemble into a feature vector $f_{v}$. Recall that this can be viewed as using the empirical approximation $\rho_{h}^{(N)}\approx\rho_{h}$ given by (4.2).
To learn the correction terms for each particle, we then train a neural network $F^{\mathrm{gain}}:\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{\text{ST}}}\rightarrow\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}$ with the standard multilayer perceptron (MLP) architecture that takes the feature vector for the ensemble as an input:
@@
From (4.27a) and (4.27b), it is evident that if the correction terms satisfy $\widehat{w}^{(n)}_{\theta}=\widehat{z}^{(n)}_{\theta}=0$, the formulation reduces to the EnKF (2.16). When the dynamics and observation models are linear, or in other words, when the filtering distribution of the states $v$ is Gaussian, the Kalman filter achieves optimality [7, 8, 40]. Therefore, the motivation for our proposed scheme (2.22) lies in its optimality in the linear (or Gaussian) setting, while further aiming to improve the performance of the Kalman filter in nonlinear problems by learning the correction terms.
Remark 9 (On the Invertibility in the Gain Computation).
The matrix $K_{\theta}^{(2)}$, defined in (4.27b) as an average of outer products, is positive semi-definite. Since $\Gamma$ is a fixed positive definite matrix, the sum $K_{\theta}^{(2)}+\Gamma$ is also positive definite and thus invertible. Crucially, the minimum eigenvalue of $K_{\theta}^{(2)}+\Gamma$ is bounded from below by the minimum eigenvalue of $\Gamma$, which is a positive constant independent of any trainable parameters. This property ensures the numerical stability of computing $(K_{\theta}^{(2)}+\Gamma)^{-1}$ during backpropagation.
@@
In addition to the proposed approach with correction terms, we have implemented the end-to-end learning of the operator $\mathfrak{B}$ presented in Subsection 2.4 and learning the Kalman gain $K_{\theta}$ directly without any constraint on its form. We find that the end-to-end approach or directly learning $K_{\theta}$ results in unstable and slower learning. A poorly initialized or an unconstrained $K_{\theta}$ leads to substantial divergence of the resulting filtered trajectory away from the ground truth trajectory. The EnKF-like structure in (4.26b) with learnable correction terms provides better stability while maintaining adaptability to nonlinear dynamics.
@@ 4.2.2 Learning the Inflation
The most common form of inflation is multiplicative inflation, where the analysis ensemble $\{v^{(n)}\}_{n=1}^{N}$ that is calculated after the analysis step (4.1c) is modified as follows
@@
where $\alpha\geq 1$ is the inflation parameter. In our learning framework, we replace the term proportional to $(\alpha-1)$ by a learned correction term $\widehat{u}^{(n)}_{\theta}$,
that is output by an MLP-based neural network $F^{\mathrm{infl}}:\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{\text{ST}}}\rightarrow\mathbb{R}^{d_{v}}$ given by
$F^{\mathrm{infl}}\bigl{(}v^{(n)},f_{v};\theta_{\mathrm{infl}}\bigr{)}=\widehat{u}^{(n)}_{\theta},$ (4.30)
where $v^{(n)}$ is the estimated state given by (4.1c), and $\theta_{\mathrm{infl}}$ denotes trainable parameters. The learned inflation step is processed after the analysis step (4.1c), resulting in the particle update:
$v^{(n)}\leftarrow v^{(n)}+\widehat{u}^{(n)}_{\theta},$ (4.31)
@@
The common form of inflation is a function of the ensemble states, i.e., not the observations, in (4.29). Similarly, the learned inflation term $\widehat{u}^{(n)}_{\theta}$ in (4.30) depends on the ensemble states $\widehat{v}^{(n)}$ and the feature vector $f_{v}$, without information of the true observation $y^{\dagger}$ or the observation noise covariance $\Gamma$. We avoid the risk that the inflation term would evolve into an end-to-end correction term, ensuring the learning performed by other components in our framework remains meaningful and effective.